A gas has a temperature of 150 K at a volume of .150 L. If the final temperature is 25 K, what is the final volume (L)?
LAW:
1
Expert's answer
2020-05-05T03:37:23-0400

pV=nRT.

n=pV/RT= 101.325 x 150 / 8.314 x 150= 12,1872745.

V= nRT/p= 12,1872745 x 8.314 x 25 / 101.325 = 25.